# Accretion flow around Kerr metric in the infra-red limit of asymptotically safe gravity

### v1

@@ Abstract
We investigate accretion disk dynamics and the formation of quasi-periodic oscillations (QPOs) in the infrared limit around Kerr-like black holes in asymptotically safe gravity. Relativistic hydrodynamic solutions of Bondi-Hoyle-Lyttleton (BHL) accretion reveal that quantum corrections significantly modify the structure of the shock cone formed around the black hole. The black hole spin controls the asymmetric of the shock cone through frame-dragging effects, whereas the quantum correction parameter softens the effective gravitational potential, resulting in a wider shock opening angle, weaker post-shock compression, and reduced density concentration within the cone. Time-dependent mass accretion rates reveal oscillation modes trapped within the shock cone. The power spectral density (PSD) investigations suggest that these modes naturally generate low-frequency QPOs, whose amplitudes, coherence, and harmonic structure depend on both the spin and the quantum correction parameter. The PSD analyses performed at different radial locations reveal that identical QPO frequencies are obtained in all cases. The numerically detected frequencies result from the excitation of global oscillation modes trapped within the post-shock region. The resulting global modes are found to consist of fundamental frequencies, their associated harmonic overtones, and near-commensurate frequency ratios such as 2:1 and 3:2. Coherent oscillations are enhanced and near-commensurate frequency ratios are produced when moderate rotation and moderate quantum corrections are coupled. Large quantum correction parameters, on the other hand, wash out unique spectral peaks and suppress oscillation amplitudes.
Keywords: Circular orbits; Particle motion; numerical simulations; shock cones; QPOs
I Introduction
@@
The possible phenomenological implications of asymptotically safe gravity are an essential feature. Compact astrophysical objects, in particular, offer a perfect testing ground for quantum gravity. Among these objects, BHs are of special interest due to the extreme curvature regimes they exhibit, as BHs represent regions of spacetime where the gravitational field becomes so intense that no signal or information, including light, can escape, rendering them inaccessible to direct observation. The event horizon delineates the ultimate limit beyond which escape is impossible. In classical BH solutions, a curvature singularity is at the core and is shielded by the event horizon; however, several theoretical constructions propose regular BH models in which singular behaviour is avoided near or at the horizon. The temporal evolution of quantum entanglement in eternal BHs embedded in anti de Sitter spacetime, along with the dependence of entropy on interior geometric properties, was analyzed in Hartman and Maldacena (2013). In addition, Engelhardt and Wall Engelhardt and Wall (2015) proposed quantum extremal surfaces, which provide a method for calculating holographic entanglement entropy that accounts for quantum corrections beyond the traditional GTR description. A variety of BH solutions incorporating asymptotic safety effects were explored in previous studies Cai and Easson (2010); Zhang (2018); Pawlowski and Tränkle (2024). From a quantum point of view, BHs are not entirely black. Hawking proved that quantum fluctuations near the event horizon produce thermal radiation Hawking (1978), which causes a slow loss of mass and a rise in temperature. This has significant effects on the thermodynamic behaviour of small black holes Gour and Medved (2003). The related thermodynamic corrections and quantum effects have also been investigated for higher-dimensional, charged, and rotating BH configurations  Pourhassan et al. (2017); Javed et al. (2023); Yousaf et al. (2024).
@@
The study of such systems relies on complementary observational techniques: spectroscopic analyses, which focus on the energy and frequency distribution of emitted photons, and timing analyses, which examine temporal variations in photon flux. Together, these methods form a robust framework for investigating accretion dynamics and strong-field gravity effects Remillard and McClintock (2006a). In a related theoretical development, Ashraf et al. Ashraf et al. (2025b) examined the dynamics of test particles orbiting a Schwarzschild BH embedded in a Dehnen-type dark matter halo, demonstrating how variations in the BH mass, halo radius, and central density influence particle motion. Closed-form expressions were derived for the particle energy, angular momentum, effective potential, and ISCO radius. The authors further explored epicyclic oscillations near the equatorial plane by analyzing radial, vertical, and orbital frequencies, as well as periastron precession. Additionally, high-energy particle collisions near the event horizon were investigated through the evaluation of centre-of-mass energy, revealing a pronounced sensitivity of particle dynamics to the underlying model parameters. Complementary analyses of orbital motion and QPO diagnostics around rotating hairy BHs in Horndeski gravity were presented in Ashraf et al. (2025a). Overall, extensive research has been devoted to understanding QPO phenomena in the strong-gravity regime surrounding BHs, underscoring their significance as powerful observational probes of relativistic astrophysics Rezzolla (2004); Belloni et al. (2012).
@@
Building upon this body of work, the present study focuses on rotating BH geometries, emphasizing the role of a scale-dependent gravitational coupling motivated by the infrared behaviour of the theory. Earlier investigations of quantum-modified Kerr spacetimes were carried out in Reuter and Tuiran (2011), where the classical Newtonian constant was replaced by a radially varying function induced by quantum gravitational effects. That analysis examined modifications to the horizon structure, ergoregions, static limit surfaces, and energy-extraction processes. In contrast, the present study employs an infrared-inspired functional form of the gravitational coupling within the asymptotic safety paradigm, enabling a more concrete and systematic analysis. This gravitational background allows the derivation of explicit analytic expressions for equatorial geodesics and facilitates a reexamination of the Penrose energy-extraction mechanism in the presence of quantum-gravitational corrections. Furthermore, we investigate accretion disk behavior and the emergence of QPOs around Kerr BHs in the infrared regime of asymptotically safe gravity. Our relativistic hydrodynamic analysis of BHL accretion demonstrates that quantum effects play a crucial role in reshaping the shock cone structure formed in the vicinity of the BH, whereas frame-dragging associated with BH rotation governs the asymmetry of the cone, the quantum correction parameter effectively weakens the gravitational potential, leading to an increased opening angle, reduced post-shock compression, and a lower matter density within the shocked region. The temporal evolution of the mass accretion rate exhibits characteristic oscillatory modes confined to the shock cone; however, PSD analyses reveal that these modes give rise to LFQPOs, with their strength, coherence, and harmonic content sensitive to both the BH spin and the magnitude of the quantum corrections. Notably, intermediate values of rotation and quantum modification favour the development of coherent oscillations with nearly commensurate frequency ratios, whereas stronger quantum effects damp the oscillations and smear out well-defined spectral features.
@@
The structure of our manuscript is as follows: In Sec . II, we introduce the Kerr metric in the infrared limit of asymptotically safe gravity, while Sec . III is devoted to the analysis of accretion dynamics and shock cone formation in the infrared-modified Kerr spacetime. In Sec . IV, we examine the impact of quantum gravitational corrections on QPOs through a detailed PSD analysis. Furthermore, Sec. V focuses on the identification of observationally relevant QPO frequencies extracted from the numerically obtained PSDs.
II Kerr metric in the infra-red limit of asymptotically safe gravity theory
@@
where $\Sigma=r^{2}+a^{2}\cos^{2}\theta$ and $\Delta=r^{2}-2Mr+\frac{2M\xi}{r}+a^{2}$. This metric reduces to its static, spherically symmetric form when $a\rightarrow 0$. To aid the reader’s understanding and provide stronger grounds for the results computed in the rest of the sections, a detailed derivation of the metric in Eq. (II) using the technique in Azreg-Aïnou (2014) is presented in the appendix.
@@
Figure 5: Numerical results for the mass accretion properties of the quantum-corrected Kerr black hole in the infra-red limit of asymptotically safe gravity. The figure is the same as Fig.4, but in this case the mass accretion rate $dM/dt$, the mean accretion rate $\langle\dot{M}\rangle$, and the fractional RMS variability are computed at $r=6.11M$, corresponding to the vicinity of the ISCO.
Figure 6: The PSD analysis is performed for both the Schwarzschild black hole and quantum-corrected black holes with different values of the parameter $\xi$, using the temporal oscillations of the mass accretion rate after it reaches saturation around a nonrotating black hole. The left panel shows the PSD computed from the mass accretion rate extracted at $r=2.3M$, corresponding to the region where matter falls toward the black hole, while the right panel presents the PSD evaluated near the ISCO at $r=6.11M$. The figures provide a comparative illustration of the variations in the QPO characteristics as a function of the black hole spin parameter $a$ and the quantum-correction parameter $\xi$. All frequencies are scaled to a black hole mass of $M=10\,M_{\odot}$.
@@
III Accretion Dynamics and Shock Phenomenology in Infra-Red Modified Kerr Geometry
In this section, we investigate the accretion dynamics forming around a black hole described by the Kerr metric in the infrared limit of asymptotically safe gravity. In order to achieve this, we numerically solve the general relativistic hydrodynamic (GRHD) equations with the spacetime metric given in Eq.II. The evolution of the fluid is carried out using a high-resolution shock-capturing (HRSC) scheme, which enables an accurate and stable treatment of nonlinear relativistic flows as well as the formation of strong discontinuities such as shocks Donmez (2004, 2006). The hydrodynamic equations are written in conservative form and are evolved on a fixed quantum-corrected Kerr spacetime background. As a result of these calculations, deviations from the classical Kerr solution induced by the quantum correction term are clearly revealed in the numerical results.
@@ III Accretion Dynamics and Shock Phenomenology in Infra-Red Modified Kerr Geometry
In relativistic settings, particularly in rotating black hole spacetimes, the structure and stability of the shock cone are influenced by frame-dragging effects and the detailed form of the gravitational potential. Consequently, physical properties such as the shock cone’s morphology, opening angle, and the density and velocity distributions of the matter confined within the cone play a crucial role in testing the Kerr black hole in the infrared limit of asymptotically safe gravity and in revealing its deviations from the classical Kerr solution.
@@
In the top row of Fig. 2, the effect of the quantum correction parameter $\xi$ is demonstrated for a non-rotating black hole ($a=0M$) by comparing shock cone structures obtained for different values of $\xi$. In the left panel, corresponding to the IR1 model, the spacetime admits both inner and outer horizons, and the accretion flow develops a well-defined and narrow shock cone in the downstream region. Due to strong collimation of matter, a dense post-shock region is formed, and the velocity vectors indicate a smooth gravitational focusing of the inflow toward the black hole. On the otherhand, for the IR2 model with the same spin parameter but a larger value of $\xi$, we have a naked singularity spacetime . In this case, the resulting shock cone is broader and exhibits weaker collimation of matter. The density contrast along the shock is reduced, and the velocity field indicates a less efficient focusing of the inflowing material. It demonstrates that increasing the quantum correction parameter $\xi$ in non-rotating configurations effectively softens the gravitational potential, thereby weakening shock compression and modifying the accretion flow.
The middle row of Fig. 2 shows the IR3 and IR4 models for a moderately rotating black hole with spin parameter $a=0.5M$, illustrating the combined influence of rotation and quantum corrections. As is well known, rotation leads to a highly asymmetric shock cone in the downstream region due to frame-dragging. When comparing IR3 and IR4 models, we find that for the larger value of $\xi$ (IR4), the shock opening angle widens and the cone structure is less sharply defined. Although frame-dragging remains present, its impact on the shock geometry is partially mitigated by the quantum correction. As $\xi$ increases, the quantity of stuff trapped inside the shock cone drops and the contour lines grow smoother, suggesting a less abrupt transition across the shock. Larger values of $\xi$ thereby lessen strong gravitational focusing and the relative impact of the spin parameter on the shock structure, even in rotating black hole spacetimes.
Finally, the bottom row of Fig. 2 presents models IR5 and IR6 corresponding to a rapidly rotating black hole with $a=0.9M$, where the effects of both strong rotation and the quantum correction parameter $\xi$ are most pronounced. The shock cone is severely bent and warped by the strong gravitational field induced by rapid rotation, creating a highly asymmetric shape. A comparatively greater shock opening angle and a less elongated shock cone result from the partial weakening of the spin parameter’s influence as $\xi$ grows. As a result, the density of stuff accumulated inside the cone drops. The strong curvature effects characteristic of rapidly rotating Kerr black holes are counteracted by the quantum correction parameter, even if rotation still imprints asymmetry on the flow.
@@
A comparison between the Schwarzschild and Kerr solutions further shows that the black hole spin parameter $a$ induces a systematic azimuthal shift in the location of the maximum density within the shock cone. A redistribution of the post-shock material accompanies this shift and arises directly from frame dragging. On other hand, density profiles of Kerr black holes with asymptotically safe gravity are consistently less peaked and more stretched than those of classical Kerr black holes with the same spin. The modified metric functions, namely the horizon function $\Delta=r^{2}-2Mr+\frac{2M\xi}{r}+a^{2}$, and the factors $(1-\xi/r^{2})$ that show up in the components $g_{tt}$, $g_{t\phi}$, and $g_{\phi\phi}$, are the source of these deviations. These parameters significantly reduce the gravitational potential in the strong-field region, which lowers the gravitational focusing efficiency. In the strong-field region, these terms effectively weaken the gravitational potential, thereby reducing the efficiency of gravitational focusing. As a result, the shock compression becomes less effective, the densest part of the post-shock flow spreads over a wider azimuthal region, and the maximum rest-mass density decreases. This behaviour is consistently observed for all considered spin configurations when different values of the quantum-correction parameter $\xi$ are taken into account. of the quantum-correction parameter $\xi$ are taken into account.
@@
In Fig.4, the top panel shows the time evolution of the mass accretion rate $dM/dt$ measured at $r=2.3M$, which is the closest extraction radius to the compact object considered in our simulations. The accretion rate is evaluated only after the system has reached a quasi-steady state, namely from $t\simeq 3000M$ up to the end of the numerical integration. In this regime, the shock cone has been formed and settled, and the residual temporal variations of $dM/dt$ represent physical oscillations of the inner accretion flow. In general, the oscillatory behavior of the numerically obtained mass accretion rate is governed by three main physical mechanisms: (i) the strength of gravitational focusing, which determines how efficiently upstream matter is concentrated into the downstream shock cone, (ii) the degree of post-shock compression inside the cone, which controls the density enhancement and pressure gradients in the inner flow, and (iii) frame dragging effects induced by black hole rotation, which bend streamlines, modify angular momentum transport, and regulate the stability of the shock structure. In the Kerr black hole of asymptotically safe gravity, the lapse and dragging factors differ from the Kerr black hole via a factor of the form $(1-\xi/r^{2})$, which alters the structure of spacetime in the strong-field region and, consequently, changes the dynamical boundary conditions experienced by the accreting fluid. Such effects remain important even in the $a=0$, $\xi=0.6M^{2}$ configuration, where no classical event horizon is present.
@@
In Fig.5, the same physical quantities discussed in detail for Fig.4 are examined at $r=6.11M$, corresponding to a location slightly farther from the strong gravity region and close to the ISCO. The temporal history of the mass accretion rate is still generally consistent among models when compared to Fig. 4, but the oscillation amplitudes are more prominent and, in some cases, amplified in comparison to those seen at smaller radii. This suggests that large-scale flow modulation and shock-cone breathing modes remain dynamically significant near the ISCO.
@@
These results agree with theoretical expectations: quantum-corrected gravity effects predominantly modify the innermost strong-field region near the black hole, whereas at ISCO the accretion variability is increasingly governed by shock-cone dynamics and larger-scale flow structures. Consequently, the modulation induced by the shock cone becomes more prominent at $r=6.11M$. In contrast, the direct imprint of quantum corrections on the accretion flow is reduced compared to the behaviour observed closer to the compact object.
IV Quantum-Correction Effects on QPOs: Power Spectral Density Analysis

### v2

@@ Abstract
We investigate accretion disk dynamics and the formation of quasi-periodic oscillations (QPOs) in the infrared limit around Kerr-like black holes in asymptotically safe gravity. Relativistic hydrodynamic solutions of Bondi-Hoyle-Lyttleton (BHL) accretion reveal that quantum corrections significantly modify the structure of the shock cone formed around the black hole. The black hole spin controls the azimuthal asymmetry of the shock cone through frame-dragging effects, whereas the quantum correction parameter effectively reduces the strength of gravitational focusing by modifying the metric coefficients in the strong-field region, resulting in a wider shock opening angle, weaker post-shock compression, and reduced density concentration within the cone. Time-dependent mass accretion rates reveal oscillation modes trapped within the shock cone. The power spectral density (PSD) investigations suggest that these modes naturally generate low-frequency QPOs, whose amplitudes, coherence, and harmonic structure depend on both the spin and the quantum correction parameter. The PSD analyses performed at different radial locations reveal that identical QPO frequencies are obtained in all cases. The numerically detected frequencies result from the excitation of global oscillation modes trapped within the post-shock region. The resulting global modes are found to consist of fundamental frequencies, their associated harmonic overtones, and near-commensurate frequency ratios such as 2:1 and 3:2. Coherent oscillations are enhanced and near-commensurate frequency ratios are produced when moderate rotation and moderate quantum corrections are coupled. Large quantum correction parameters, on the other hand, wash out unique spectral peaks and suppress oscillation amplitudes.
Keywords:
Circular orbits, numerical simulations, shock cones, QPOs, XRBs
1 Introduction
@@
The possible phenomenological implications of asymptotically safe gravity are an essential feature. Compact astrophysical objects, in particular, offer a perfect testing ground for quantum gravity. Among these objects, BHs are of special interest due to the extreme curvature regimes they exhibit, as BHs represent regions of spacetime where the gravitational field becomes so intense that no signal or information, including light, can escape, rendering them inaccessible to direct observation. The event horizon delineates the ultimate limit beyond which escape is impossible. In classical BH solutions, a curvature singularity is at the core and is shielded by the event horizon; however, Several theoretical constructions propose regular BH models in which the central curvature singularity is resolved or avoided, while the event horizon itself remains classically regular. The temporal evolution of quantum entanglement in eternal BHs embedded in anti de Sitter spacetime, along with the dependence of entropy on interior geometric properties, was analyzed in Hartman and Maldacena (2013). In addition, Engelhardt and Wall Engelhardt and Wall (2015) proposed quantum extremal surfaces, which provide a method for calculating holographic entanglement entropy that accounts for quantum corrections beyond the traditional GTR description. A variety of BH solutions incorporating asymptotic safety effects were explored in previous studies Cai and Easson (2010); Zhang (2018); Pawlowski and Tränkle (2024). From a quantum point of view, BHs are not entirely black. Hawking proved that quantum fluctuations near the event horizon produce thermal radiation Hawking (1978), which causes a slow loss of mass and a rise in temperature. This has significant effects on the thermodynamic behaviour of small black holes Gour and Medved (2003). The related thermodynamic corrections and quantum effects have also been investigated for higher-dimensional, charged, and rotating BH configurations  Pourhassan et al. (2017); Javed et al. (2023); Yousaf et al. (2024).
@@
The study of such systems relies on complementary observational techniques: spectroscopic analyses, which focus on the energy and frequency distribution of emitted photons, and timing analyses, which examine temporal variations in photon flux. Together, these methods form a robust framework for investigating accretion dynamics and strong-field gravity effects Remillard and McClintock (2006a). In a related theoretical development, Ashraf et al. Ashraf et al. (2025b) examined the dynamics of test particles orbiting a Schwarzschild BH embedded in a Dehnen-type dark matter halo, demonstrating how variations in the BH mass, halo radius, and central density influence particle motion. Closed-form expressions were derived for the particle energy, angular momentum, and effective potential. The authors further explored epicyclic oscillations near the equatorial plane by analyzing radial, vertical, and orbital frequencies, as well as periastron precession. Additionally, high-energy particle collisions near the event horizon were investigated through the evaluation of centre-of-mass energy, revealing a pronounced sensitivity of particle dynamics to the underlying model parameters. Complementary analyses of orbital motion and QPO diagnostics around rotating hairy BHs in Horndeski gravity were presented in Ashraf et al. (2025a). Overall, extensive research has been devoted to understanding QPO phenomena in the strong-gravity regime surrounding BHs, underscoring their significance as powerful observational probes of relativistic astrophysics Rezzolla (2004); Belloni et al. (2012).
@@
Building upon this body of work, the present study focuses on rotating black hole geometries described by an infrared-effective metric arising from the renormalization-group flow of asymptotically safe gravity, characterized by a scale-dependent gravitational coupling. Earlier investigations of quantum-modified Kerr spacetimes were carried out in Reuter and Tuiran (2011), where the classical Newtonian constant was replaced by a radially varying function induced by quantum gravitational effects. That analysis examined modifications to the horizon structure, ergoregions, static limit surfaces, and energy-extraction processes. In contrast, the present study employs an infrared-inspired functional form of the gravitational coupling within the asymptotic safety paradigm, enabling a more concrete and systematic analysis. This gravitational background allows the derivation of explicit analytic expressions for equatorial geodesics and facilitates a reexamination of the Penrose energy-extraction mechanism in the presence of quantum-gravitational corrections. Furthermore, we investigate accretion disk behavior and the emergence of QPOs around Kerr BHs in the infrared regime of asymptotically safe gravity. Our relativistic hydrodynamic analysis of BHL accretion demonstrates that quantum effects play a crucial role in reshaping the shock cone structure formed in the vicinity of the BH, whereas frame-dragging associated with BH rotation governs the asymmetry of the cone, the quantum correction parameter effectively weakens the gravitational potential in this relativistic Kerr-like spacetime, leading to an increased opening angle, reduced post-shock compression, and a lower matter density within the shocked region. The temporal evolution of the mass accretion rate exhibits characteristic oscillatory modes confined to the shock cone; however, PSD analyses reveal that these modes give rise to LFQPOs, with their strength, coherence, and harmonic content sensitive to both the BH spin and the magnitude of the quantum corrections. Notably, intermediate values of rotation and quantum modification favour the development of coherent oscillations with nearly commensurate frequency ratios, whereas stronger quantum effects damp the oscillations and smear out well-defined spectral features.
@@
The structure of our manuscript is as follows: In Sec . 2, we introduce the Kerr metric in the infrared limit of asymptotically safe gravity, while Sec . 3 is devoted to the analysis of accretion dynamics and shock cone formation in the infrared-modified Kerr spacetime. In Sec . 4, we examine the impact of quantum gravitational corrections on QPOs through a detailed PSD analysis. Furthermore, Sec. 5 focuses on the identification of observationally relevant QPO frequencies extracted from the numerically obtained PSDs.
2 Kerr metric in the infra-red limit of asymptotically safe gravity theory
@@
where $\Sigma=r^{2}+a^{2}\cos^{2}\theta$ and $\Delta=r^{2}-2Mr+\frac{2M\xi}{r}+a^{2}$. This metric reduces to its static, spherically symmetric form when $a\rightarrow 0$. To aid the reader’s understanding and provide stronger grounds for the results computed in the rest of the sections, a detailed derivation of the metric in Eq. (2) using the technique in Azreg-Aïnou (2014) is presented in the appendix.
@@
Figure 5: Numerical results for the mass accretion properties of the quantum-corrected Kerr black hole in the infra-red limit of asymptotically safe gravity. The figure is the same as Fig.4, but in this case the mass accretion rate $dM/dt$, the mean accretion rate $\langle\dot{M}\rangle$, and the fractional RMS variability are computed at $r=6.11M$, corresponding to a region farther from the immediate vicinity of the black hole.
Figure 6: The PSD analysis is performed for both the Schwarzschild black hole and quantum-corrected black holes with different values of the parameter $\xi$, using the temporal oscillations of the mass accretion rate after it reaches saturation around a nonrotating black hole. The left panel shows the PSD computed from the mass accretion rate extracted at $r=2.3M$, corresponding to the region where matter falls toward the black hole, while the right panel presents the PSD evaluated at $r=6.11M$. The figures provide a comparative illustration of the variations in the QPO characteristics as a function of the black hole spin parameter $a$ and the quantum-correction parameter $\xi$. All frequencies are scaled to a black hole mass of $M=10\,M_{\odot}$.
@@
3 Accretion Dynamics and Shock Phenomenology in Infra-Red Modified Kerr Geometry
In this section, we investigate the accretion dynamics forming around a black hole described by the Kerr metric in the infrared limit of asymptotically safe gravity. In order to achieve this, we numerically solve the general relativistic hydrodynamic (GRHD) equations with the spacetime metric given in Eq.2. The evolution of the fluid is carried out using a high-resolution shock-capturing (HRSC) scheme, which enables an accurate and stable treatment of nonlinear relativistic flows as well as the formation of strong discontinuities such as shocks Donmez (2004, 2006). The hydrodynamic equations are written in conservative form and are evolved on a fixed quantum-corrected Kerr spacetime background. As a result of these calculations, deviations from the classical Kerr solution induced by the quantum correction term are clearly revealed in the numerical results.
@@ 3 Accretion Dynamics and Shock Phenomenology in Infra-Red Modified Kerr Geometry
In relativistic settings, particularly in rotating black hole spacetimes, the structure and stability of the shock cone are influenced by frame-dragging effects and the detailed form of the gravitational attraction. Consequently, physical properties such as the shock cone’s morphology, opening angle, and the density and velocity distributions of the matter confined within the cone play a crucial role in testing the Kerr black hole in the infrared limit of asymptotically safe gravity and in revealing its deviations from the classical Kerr solution.
@@
In the top row of Fig. 2, the effect of the quantum correction parameter $\xi$ is demonstrated for a non-rotating black hole ($a=0M$) by comparing shock cone structures obtained for different values of $\xi$. In the left panel, corresponding to the IR1 model, the spacetime admits both inner and outer horizons, and the accretion flow develops a well-defined and narrow shock cone in the downstream region. Due to strong collimation of matter, a dense post-shock region is formed, and the velocity vectors indicate a smooth gravitational focusing of the inflow toward the black hole. On the other hand, for the IR2 model with the same spin parameter but a larger value of $\xi$, we have a naked singularity spacetime . In this case, the resulting shock cone is broader and exhibits weaker collimation of matter. The density contrast along the shock is reduced, and the velocity field indicates a less efficient focusing of the inflowing material. It demonstrates that increasing the quantum correction parameter $\xi$ in non-rotating configurations effectively softens the gravitational attraction, thereby weakening shock compression and modifying the accretion flow.
The middle row of Fig. 2 shows the IR3 and IR4 models for a moderately rotating black hole with spin parameter $a=0.5M$, illustrating the combined influence of rotation and quantum corrections. As is well known, black hole rotation induces a pronounced azimuthal asymmetry of the downstream shock cone through frame-dragging effects. When comparing the IR3 and IR4 models, we find that for the larger value of $\xi$ (IR4), the shock opening angle widens and the cone structure is less sharply defined. Although frame-dragging remains present, its impact on the shock geometry is partially mitigated by the quantum correction. As $\xi$ increases, the quantity of stuff trapped inside the shock cone drops and the contour lines grow smoother, suggesting a less abrupt transition across the shock. Larger values of $\xi$ thereby lessen strong gravitational focusing and the relative impact of the spin parameter on the shock structure, even in rotating black hole spacetimes.
Finally, the bottom row of Fig. 2 presents models IR5 and IR6 corresponding to a rapidly rotating black hole with $a=0.9M$, where the effects of both strong rotation and the quantum correction parameter $\xi$ are most pronounced. The shock cone is strongly bent and azimuthally distorted by frame-dragging associated with rapid rotation, resulting in a pronounced azimuthal asymmetry. A comparatively greater shock opening angle and a less elongated shock cone result from the partial weakening of the spin parameter’s influence as $\xi$ grows. As a result, the density of stuff accumulated inside the cone drops. The strong curvature effects characteristic of rapidly rotating Kerr black holes are counteracted by the quantum correction parameter, even if rotation still imprints asymmetry on the flow.
@@
A comparison between the Schwarzschild and Kerr solutions further shows that the black hole spin parameter $a$ induces a systematic azimuthal shift in the location of the maximum density within the shock cone. A redistribution of the post-shock material accompanies this shift and arises directly from frame dragging. On other hand, density profiles of Kerr black holes with asymptotically safe gravity are consistently less peaked and more stretched than those of classical Kerr black holes with the same spin. The modified metric functions, namely the horizon function $\Delta=r^{2}-2Mr+\frac{2M\xi}{r}+a^{2}$, and the factors $(1-\xi/r^{2})$ that show up in the components $g_{tt}$, $g_{t\phi}$, and $g_{\phi\phi}$, are the source of these deviations. These parameters significantly reduce the gravitational potential in the strong-field region of the relativistic Kerr-like spacetime, which lowers the gravitational focusing efficiency. In the strong-field region, these terms effectively weaken the gravitational attraction, thereby reducing the efficiency of gravitational focusing. As a result, the shock compression becomes less effective, the densest part of the post-shock flow spreads over a wider azimuthal region, and the maximum rest-mass density decreases. This behaviour is consistently observed for all considered spin configurations when different values of the quantum-correction parameter $\xi$ are taken into account. of the quantum-correction parameter $\xi$ are taken into account.
@@
In Fig.4, the top panel shows the time evolution of the mass accretion rate $dM/dt$ measured at $r=2.3M$, which is the closest extraction radius to the compact object considered in our simulations. The accretion rate is evaluated only after the system has reached a quasi-steady state, namely from $t\simeq 3000M$ up to the end of the numerical integration. In this regime, the shock cone has been formed and settled, and the residual temporal variations of $dM/dt$ represent physical oscillations of the inner accretion flow. In general, the oscillatory behavior of the numerically obtained mass accretion rate is governed by three main physical mechanisms: (i) the strength of gravitational focusing, which determines how efficiently upstream matter is concentrated into the downstream shock cone, (ii) the degree of post-shock compression inside the cone, which controls the density enhancement and pressure gradients in the inner flow, and (iii) frame dragging effects induced by black hole rotation, which bend streamlines, modify angular momentum transport, and regulate the stability of the shock structure. In the Kerr black hole of asymptotically safe gravity, the lapse and dragging factors differ from the Kerr black hole via a factor of the form $(1-\xi/r^{2})$, which alters the structure of spacetime in the strong-field region and, consequently, changes the dynamical boundary conditions experienced by the accreting fluid. Such effects remain important even in the $a=0$, $\xi=0.6M^{2}$ configuration, which corresponds to a horizonless spacetime, but in this case, the GRHD evolution is performed outside an excised inner region, where the metric remains regular and the fluid dynamics are well defined.
@@
In Fig.5, the same physical quantities discussed in detail for Fig.4 are examined at $r=6.11M$, corresponding to a location slightly farther from the strong gravity region. The temporal history of the mass accretion rate is still generally consistent among models when compared to Fig. 4, but the oscillation amplitudes are more prominent and, in some cases, amplified in comparison to those seen at smaller radii. This suggests that large-scale flow modulation and shock-cone breathing modes remain dynamically significant away from the immediate vicinity of the black hole.
@@
These results agree with theoretical expectations: quantum-corrected gravity effects predominantly modify the innermost strong-field region near the black hole, whereas at larger radii the accretion variability is increasingly governed by shock-cone dynamics and larger-scale flow structures. Consequently, the modulation induced by the shock cone becomes more prominent at $r=6.11M$. In contrast, the direct imprint of quantum corrections on the accretion flow is reduced compared to the behaviour observed closer to the compact object.
Overall, Since both the quantum-correction parameter $\xi$ and the black hole spin parameter $a$ significantly influence the dynamical structure of the shock cone and the associated variability properties, their impact on the resulting physical processes is qualitatively substantial. The black hole spin parameter primarily controls the azimuthal displacement of the shock cone and induces azimuthal asymmetry through frame-dragging effects. In contrast, the quantum-correction parameter $\xi$ symmetrically modifies the strength of gravitational focusing and the compression of matter within the shock cone. As a result, variations in $\xi$ lead to changes in the shock opening angle, post-shock density contrast, and the efficiency with which perturbations are trapped within the post-shock cavity—effects that cannot be reproduced by variations in the spin parameter alone.
4 Quantum-Correction Effects on QPOs: Power Spectral Density Analysis
In the present simulations, the global oscillation modes arise from the presence of a well-defined shock cone formed in the downstream region of the black hole. The shock cone creates a finite post-shock cavity that acts as a resonant region in which perturbations can be trapped and sustained. Radially, these trapped modes oscillate between the inner boundary of the computational domain, where the shock cone is attached, and the stagnation region of the flow, which is typically located around $r\simeq 26M$ along the downstream axis ($\phi=0$). At the same time, the flow can oscillate laterally between the boundaries of the shock cone in the azimuthal direction. As a result, pressure and velocity perturbations are confined within the entire post-shock cavity and undergo repeated radial and azimuthal reflections. The characteristic timescales of these oscillations are therefore set by the advection and sound-crossing times across the cavity, leading to coherent global oscillation modes rather than localized, radius-dependent fluctuations. Further details on these modes and their physical origin can be found in Refs. Cruz-Osorio et al. (2023); Donmez (2024).